Are nursing salaries in Tampa, Florida, lower than those in Dallas, Texas? As reported by the Tampa Tribune, salary data show staff nurses in Tampa earn less than staff nurses in Dallas. Suppose that in a follow-up study of 35 staff nurses in Tampa and 45 staff nurses in Dallas you obtain the following results. Enter negative value as negative number. ● Tampa Dallas m, = 35 1 $54,100 81 = $4,800 i2 . $56,600 $5,500 a. Formulate hypothesis so that, if the null hypothesis is rejected, we can conclude that salaries for staff nurses in Tampa are significantly lower than for those i Dallas. Use α-0.05. Ha : μί-μ2 C-Select your answer-3 0 . : μι-μ2 C select your answer-) 0 b. What is the value of the test statistic? (to 3 decimals) c. What is the p-value? Degrees of freedom isRound degrees of freedom to previous whole number p-value is Select your answer3 d. What is your conclusion? We - Select your answer 3 that the salaries of staff nurses are lower in Tampa than in Dallas
Answer #1

a)

H_0:\mu _1-\mu _2=0

H_a:\mu _1-\mu _2<0

b)standard error for difference=sqrt((4800^2/35)+(5500^2/45))=1153.476

Test statistic:

t=(54100-56600)/1153.476

t=-2.167

c)df=76.92

df=76

p-value=tdist(2.167,76,1)=0.0167

d)As,p-value<0.05,we reject null hypothesis.

We can conclude that the salaries of staff nurses are lower in tampa than dallas.